# Riemannian Laplace Approximation with the Fisher Metric

### v7

@@ 2.1 Laplace Approximation
Let $\operatorname{\boldsymbol{\theta}}\in\operatorname{\boldsymbol{\Theta}}$ denote the $D$-dimensional vector in the parameter space $\operatorname{\boldsymbol{\Theta}}$. Denote the log-posterior distribution given observed data set $\operatorname{\boldsymbol{y}}$ $=$ $\{y_{n}\}_{n=1}^{N}$ as $\ell_{\operatorname{\boldsymbol{y}}}(\operatorname{\boldsymbol{\theta}})$ :=assign:=:= $\log\pi(\operatorname{\boldsymbol{\theta}}|\operatorname{\boldsymbol{y}})$. Laplace’s method expands a second-order Taylor series of $\ell_{\operatorname{\boldsymbol{y}}}(\operatorname{\boldsymbol{\theta}})$ at the mode (MAP) where the first-order term disappears. This is equivalent to using a Gaussian approximation
@@
where $\hat{\operatorname{\boldsymbol{\theta}}}$ is the MAP estimate and $\nabla^{2}\ell_{\operatorname{\boldsymbol{y}}}(\hat{\operatorname{\boldsymbol{\theta}}})$ is the Hessian matrix of the log-posterior evaluated at $\hat{\operatorname{\boldsymbol{\theta}}}$. We refer to this as ELA, for Euclidean LA.
@@ 3.1 Principle and algorithm
Algorithm 1 RLA-B, producing $N$ samples $\operatorname{\boldsymbol{\theta}}^{[n]}$.
@@
Consider the special case of $g$ being the Euclidean inner product. Then $\operatorname{\boldsymbol{G}}(\operatorname{\boldsymbol{\theta}})=\operatorname{\boldsymbol{I}}_{D}$ and the exponential map becomes $\mathrm{Exp}_{\hat{\operatorname{\boldsymbol{\theta}}}}(\operatorname{\boldsymbol{v}})$ $=$ $\hat{\operatorname{\boldsymbol{\theta}}}+\operatorname{\boldsymbol{v}}$, a unit step in the direction of the velocity $\operatorname{\boldsymbol{v}}$. If velocities are sampled using $\operatorname{\boldsymbol{\Sigma}}=-(\nabla^{2}\ell_{\operatorname{\boldsymbol{y}}}(\hat{\operatorname{\boldsymbol{\theta}}}))^{-1}$ we get the classical ELA.
@@
Figure 3: Wasserstein distances from approximate samples to true samples for isotropic Gaussians of varying $D$, computed for the first dimension. The lines are means and $\text{means}\pm 2.0$ times standard deviations (stds) computed over $5$ runs.
@@ 4.1 RLA-B with logarithmic map
Consider two manifolds $\operatorname{\mathcal{P}}$ and $\operatorname{\mathcal{M}}$. $\operatorname{\mathcal{P}}=(\operatorname{\boldsymbol{\Theta}},g_{\operatorname{\mathcal{P}}})$ has the metric tensor
@@
Suppose we sample from a Gaussian distribution as in ELA and consider the samples as points on manifold $\operatorname{\mathcal{P}}$. If $\operatorname{\mathcal{P}}$ is equal to $\operatorname{\mathcal{M}}$ the transformation given by $\textrm{Exp}_{\operatorname{\mathcal{M}},\hat{\operatorname{\boldsymbol{\theta}}}}\circ\textrm{Log}_{\operatorname{\mathcal{P}},\hat{\operatorname{\boldsymbol{\theta}}}}$ results in exact samples from the target associated with $\operatorname{\mathcal{M}}$. For non-Gaussian targets the method is not exact, but we observed it can still help reducing the underestimation tendency of RLA-B.
In Algorithm 2 the exponential map is computed as before on manifold $\operatorname{\mathcal{M}}$, and the logarithmic map corresponds to a boundary value problem of the geodesic ODE on manifold $\operatorname{\mathcal{P}}$. It is solved numerically and benefits again from efficient computation of Christoffel symbols. The logarithmic map is in general more costly than the exponential map (Arvanitidis et al.,, 2016), but we observed it can help reducing the number of function evaluations during the exponential map and the overall cost is not necessarily higher. We refer to this variant as RLA-BLog.
Algorithm 2 RLA-BLog, producing $N$ samples $\operatorname{\boldsymbol{\theta}}^{[n]}$. The manifold $\operatorname{\mathcal{P}}$ uses the metric induced by Gaussian $\mathcal{N}(\tilde{\operatorname{\boldsymbol{\theta}}}|\hat{\operatorname{\boldsymbol{\theta}}},\operatorname{\boldsymbol{\Sigma}})$ and the manifold $\operatorname{\mathcal{M}}$ the metric induced by the target $\pi(\operatorname{\boldsymbol{\theta}}|\operatorname{\boldsymbol{y}})$.
@@ 4.3.2 Theoretical basis
With $(\operatorname{\boldsymbol{G}}_{\operatorname{\boldsymbol{\Psi}}})_{i,j}$ $=$ $\operatorname{\mathbb{E}}_{\operatorname{\boldsymbol{\psi}}}\big{(}-\partial^{2}_{\mu_{i},\mu_{j}}\log\mathcal{N}(\operatorname{\boldsymbol{\psi}}|\operatorname{\boldsymbol{\mu}},\operatorname{\boldsymbol{S}})\big{)}=\operatorname{\boldsymbol{S}}^{-1}_{i,j}$ inspired by the Fisher metric, we have
@@ 4.3.3 Computation of the Fisher metric
where $\boldsymbol{J}_{n}$ $=$ $[\nabla(\boldsymbol{\phi}_{n})_{1}\cdots\nabla(\boldsymbol{\phi}_{n})_{P}]^{\top}$ is a $P\times D$ Jacobian matrix of the NN at the $n^{th}$ input. In other words, we can form the FIM for the whole network by transforming the FIM on its basic form with Jacobians that can be computed using standard automatic differentiation. Formally, this is a pullback metric from the parameter space $\operatorname{\boldsymbol{\Phi}}$ to the NN parameter space $\operatorname{\boldsymbol{\Theta}}$; see Section 14.2 in the Supplement for further details.
@@ 5.1 Experimental setup
[…]
@@ 5.2 Banana distribution
[…]
@@
[…]
@@ 5.3 Bayesian logistic regression
[…]
where $\sigma:\mathcal{X}\rightarrow(0,1)$ is the sigmoid function and $\alpha=100$. The parameter space of the Bernoulli model in its basic form is $\operatorname{\boldsymbol{\phi}}\in(0,1)=\Phi$, whose FIM is known. Therefore the Fisher metric on the parameter space $\operatorname{\boldsymbol{\Theta}}$ becomes $\operatorname{\boldsymbol{G}}$ $=$ $\operatorname{\boldsymbol{X}}^{\top}\operatorname{\boldsymbol{\Lambda}}\operatorname{\boldsymbol{X}}+\alpha^{-1}\operatorname{\boldsymbol{I}}$, where $\operatorname{\boldsymbol{X}}=[\operatorname{\boldsymbol{x}}_{1}\cdots\operatorname{\boldsymbol{x}}_{N}]$ is the covariate matrix (or inputs) and $\operatorname{\boldsymbol{\Lambda}}$ is a diagonal matrix with elements $\operatorname{\boldsymbol{\Lambda}}_{nn}=\sigma\big{(}\boldsymbol{\theta}^{\top}\boldsymbol{x}_{n}\big{)}\big{(}1-\sigma(\boldsymbol{\theta}^{\top}\boldsymbol{x}_{n}\big{)}\big{)}$. We use Euclidean MAP, and here the negative Hessian coincides with the Fisher metric (Girolami and Calderhead,, 2011) (so both choices for $\operatorname{\boldsymbol{\Sigma}}$ are the same).
@@ 5.4 Neural network regression
Finally, we ran an experiment similar to the one by Bergamin et al., (2023), with data from Snelson and Ghahramani, (2005). The task is a $1$D regression problem using an NN of size $1$-$10$-$1$ with $\tanh$ activation. The methods are implemented using Daxberger et al., (2021), with Fisher precision (the default option for full LA), $\hat{\operatorname{\boldsymbol{\theta}}}$ found by standard MAP training, and post hoc optimized prior precision and noise std. MSE and NLL are computed using $500$ samples.
@@
[…]
@@ 8.1 Theorem 1
The negative Hessian is therefore constant. For the second case, where the prior is uniform, the negative Hessian is 00 everywhere, thus is also constant.
@@ 8.3 Theorem 3
there exists an isomorphism between the tangent spaces of $\operatorname{\mathcal{M}}$ and $\operatorname{\mathcal{N}}$. As a result, the exponential map on $\operatorname{\mathcal{M}}$ and $\operatorname{\mathcal{N}}$ and the tangent vectors at the center of the distribution transform correctly. Since RLA-F is exact for the distribution corresponding to $\operatorname{\mathcal{M}}$, it is also exact for the distribution corresponding to $\operatorname{\mathcal{N}}$.
@@ 9 USING HAUSDORFF MAP AND FISHER PRECISION
It is a standard result in Riemannian geometry that for a Riemannian manifold, at each tangent space there exists a normal coordinate (Lee, John M.,, 2018), such that the components of the metric at the point form an identity matrix, and all first partial derivatives of the metric vanishes at the point (see Proposition 5.24 from Lee, John M.,, 2018). Therefore, consider such a transformation to the normal coordinate $\operatorname{\boldsymbol{\phi}}:\operatorname{\boldsymbol{\psi}}\rightarrow\operatorname{\boldsymbol{\theta}}$. $\hat{\operatorname{\boldsymbol{\theta}}}=\operatorname{\boldsymbol{\phi}}(\hat{\operatorname{\boldsymbol{\psi}}})$ is the Hausdorff MAP due to Theorem 1. Recall the definition of Hausdorff MAP; it is clear that $\partialderivative{\pi^{\operatorname{\boldsymbol{G}}}(\hat{\operatorname{\boldsymbol{\theta}}})}{\theta_{i}}=0$.
@@ 11 BIAS OF RLA-B
Consider a positive $x$. It is clear that $\frac{1}{\sqrt{1+Dx^{2}}}$ decreases as $D$ increases. For smaller $D$, the expected velocity as calculated above may increase as $D$ increases, since $\operatorname{\mathbb{E}}\left[v_{0}\right]$ itself may increase. However, as $D$ approaches infinity, $\operatorname{\mathbb{E}}\left[\sqrt{D}v_{0}\right]$ approaches $\sqrt{D-1}$, $\operatorname{\mathbb{E}}\left[v_{0}\right]$ approaches $1$ and $\operatorname{\mathbb{E}}\left[v\right]$ approaches 00.
@@ 14.2 Neural Networks (NN)
where $\boldsymbol{J}_{n}$ $=$ $[\nabla(\boldsymbol{\operatorname{\boldsymbol{\phi}}}_{n})_{1}\cdots\nabla(\boldsymbol{\operatorname{\boldsymbol{\phi}}}_{n})_{P}]^{\top}$ denotes the $P\times D$ Jacobian matrix of the NN at the $n^{th}$ input.
@@ 15.3 Finding the MAP estimate
For experiments other than neural network experiments, unless there is an analytical solution of the MAP estimate, we find the MAP estimate by collecting $20$ runs of BFGS optimization with random initialization using the implementation from Virtanen et al., (2020) which yield negative log-posterior values other than plus infinity or NaN, with maximum number of iterations $61e$ 6. We choose the run that yields the smallest negative log-posterior with numpy.argmin.
For neural network experiments, we find the MAP estimate using $lr=1e-2$, $weight\_decay=1e-5$ for $20000$ epochs with the Adam (Kingma and Ba,, 2015) optimizer. The prior precision and noise std are optimized post hoc following standard procedure from Daxberger et al., (2021). During the optimization process, we fix the random seeds, in order to have consistent prior precision and noise std for NUTS.
@@ 15.5 ODE integration
For the exponential maps, other than neural network experiments, ODE integrations are carried out using Diffrax (Kidger,, 2021). Inspired by the default settings of the function solve_ivp in SciPy (Virtanen et al.,, 2020), we use Dormand-Prince’s $5(4)$ method (Dormand and Prince,, 1980). Diffrax reports the number of steps num_steps used by the integrator. In order to make the results approximately comparable to the number of function evaluations nfev reported by SciPy, we multiply num_steps by $6$, the number of function evaluations per step of Dormand-Prince’s $5(4)$ method. We use adaptive step size with initial step size chosen by the algorithm, with $rtol=1e-3$ and $atol=1e-6$, matching the default option of SciPy. We set the maximum number of steps allowed to be $4096$ (approximately equivalent to $24576$ function evaluations), following the default configuration of Diffrax. For neural network experiments, following Bergamin et al., (2023), we use SciPy’s solve_ivp and use the default hyper parameters.
@@ 15.6 Calculating the Wasserstein distance
where $\operatorname{\boldsymbol{\gamma}}$ is the transportation plan, $\operatorname{\boldsymbol{M}}$ is the distance matrix induced by the Euclidean metric for the two sets of samples, and $\operatorname{\boldsymbol{a}}$ and $\operatorname{\boldsymbol{b}}$ are uniform numbers summing up to $1$ with the lengths the same as the numbers of respective samples. The maximum number of iterations is set to be $101e$ 10.
@@ 15.10 Neural networks
In Section 5.4, we performed experiments on NN regression tasks. Similar to Bergamin et al., (2023), we use the dataset from Snelson and Ghahramani, (2005), which has $200$ data points with labels, with the $x$ coordinates distributed between 00 and $6$.
@@ 16.5.2 Standardized data
In the neural network regression experiment from the main paper, we directly run the neural network on the original dataset from Snelson and Ghahramani, (2005). We later observed that using standardized data may lead to reduced running times for all methods while reducing possible numerical issues, and we report the results in Table 5 and Figure 14. For finding the MAP estimate, we set $weight\_decay=1e-4$, with the other settings the same as for non-standardized data.
@@
[…]

### v8

@@ 2.1 Laplace Approximation
Let $\operatorname{\boldsymbol{\theta}}\in\operatorname{\boldsymbol{\Theta}}$ denote the $D$-dimensional vector in the parameter space $\operatorname{\boldsymbol{\Theta}}$. Denote the log-posterior distribution given observed data set $\operatorname{\boldsymbol{y}}$ == $\{y_{n}\}_{n=1}^{N}$ as $\ell_{\operatorname{\boldsymbol{y}}}(\operatorname{\boldsymbol{\theta}})$ :=:= $\log\pi(\operatorname{\boldsymbol{\theta}}|\operatorname{\boldsymbol{y}})$. Laplace’s method expands a second-order Taylor series of $\ell_{\operatorname{\boldsymbol{y}}}(\operatorname{\boldsymbol{\theta}})$ at the mode (MAP) where the first-order term disappears. This is equivalent to using a Gaussian approximation
@@
[…]
@@ 3.1 Principle and algorithm
Algorithm 1 RLA-B, producing NN samples $\operatorname{\boldsymbol{\theta}}^{[n]}$.
@@
Consider the special case of $g$ being the Euclidean inner product. Then $\operatorname{\boldsymbol{G}}(\operatorname{\boldsymbol{\theta}})=\operatorname{\boldsymbol{I}}_{D}$ and the exponential map becomes $\mathrm{Exp}_{\hat{\operatorname{\boldsymbol{\theta}}}}(\operatorname{\boldsymbol{v}})$ == $\hat{\operatorname{\boldsymbol{\theta}}}+\operatorname{\boldsymbol{v}}$, a unit step in the direction of the velocity $\operatorname{\boldsymbol{v}}$. If velocities are sampled using $\operatorname{\boldsymbol{\Sigma}}=-(\nabla^{2}\ell_{\operatorname{\boldsymbol{y}}}(\hat{\operatorname{\boldsymbol{\theta}}}))^{-1}$ we get the classical ELA.
@@
Figure 3: Wasserstein distances from approximate samples to true samples for isotropic Gaussians of varying $D$, computed for the first dimension. The lines are means and $\text{means}\pm 2.0$ times standard deviations (stds) computed over 55 runs.
@@ 4.1 RLA-B with logarithmic map
[…]
@@
[…]
@@ 4.3.2 Theoretical basis
With $(\operatorname{\boldsymbol{G}}_{\operatorname{\boldsymbol{\Psi}}})_{i,j}$ == $\operatorname{\mathbb{E}}_{\operatorname{\boldsymbol{\psi}}}\big(-\partial^{2}_{\mu_{i},\mu_{j}}\log\mathcal{N}(\operatorname{\boldsymbol{\psi}}|\operatorname{\boldsymbol{\mu}},\operatorname{\boldsymbol{S}})\big)=\operatorname{\boldsymbol{S}}^{-1}_{i,j}$ inspired by the Fisher metric, we have
@@ 4.3.3 Computation of the Fisher metric
where $\boldsymbol{J}_{n}$ == $\nabla(\boldsymbol{\phi}_{n})_{1}\cdots\nabla(\boldsymbol{\phi}_{n})_{P}]^{\top}$ is a $P\times D$ Jacobian matrix of the NN at the $n^{th}$ input. In other words, we can form the FIM for the whole network by transforming the FIM on its basic form with Jacobians that can be computed using standard automatic differentiation. Formally, this is a pullback metric from the parameter space $\operatorname{\boldsymbol{\Phi}}$ to the NN parameter space $\operatorname{\boldsymbol{\Theta}}$; see Section 14.2 in the Supplement for further details.
@@ 5.1 Experimental setup
Table 1: Banana distribution results as $\text{mean},\text{std}]$. Bold font indicates the best method. $\mathcal{W}$ indicates Wasserstein distance to NUTS samples while $T$ indicates the average number of function evaluations for one sample. For all evaluation metrics smaller is better.
@@ 5.2 Banana distribution
$\operatorname{\boldsymbol{\theta}}\overset{\textrm{ind.}}{\sim}\operatorname{\mathcal{N}}(0,\sigma_{\operatorname{\boldsymbol{\theta}}}^{2})\ \textrm{and}\ Y_{n}|\operatorname{\boldsymbol{\theta}}\sim\operatorname{\mathcal{N}}(\theta_{1}+\theta_{2}^{2},\sigma_{Y}^{2}),$
@@
Table 2: Logistic regression results as $\text{mean},\text{std}]$. Bold font indicates the best method. Italic font indicates the integrator reached maximum number of steps in at least one run. $\mathcal{W}$ indicates Wasserstein distance to NUTS samples while $T$ indicates the average number of function evaluations for one sample. For all evaluation metrics smaller is better.
@@ 5.3 Bayesian logistic regression
$Y_{n}|\operatorname{\boldsymbol{\theta}}\sim\textrm{Bernoulli}(\sigma(\operatorname{\boldsymbol{\theta}}^{\top}\operatorname{\boldsymbol{x}}_{n}))\ \ \textrm{and}\ \ \operatorname{\boldsymbol{\theta}}\overset{\textrm{ind.}}{\sim}\operatorname{\mathcal{N}}(0,\alpha),$
where $\sigma:\mathcal{X}\rightarrow(0,1)$ is the sigmoid function and $\alpha=100$. The parameter space of the Bernoulli model in its basic form is $\operatorname{\boldsymbol{\phi}}\in(0,1)=\Phi$, whose FIM is known. Therefore the Fisher metric on the parameter space $\operatorname{\boldsymbol{\Theta}}$ becomes $\operatorname{\boldsymbol{G}}$ == $\operatorname{\boldsymbol{X}}^{\top}\operatorname{\boldsymbol{\Lambda}}\operatorname{\boldsymbol{X}}+\alpha^{-1}\operatorname{\boldsymbol{I}}$, where $\operatorname{\boldsymbol{X}}=[\operatorname{\boldsymbol{x}}_{1}\cdots\operatorname{\boldsymbol{x}}_{N}]$ is the covariate matrix (or inputs) and $\operatorname{\boldsymbol{\Lambda}}$ is a diagonal matrix with elements $\operatorname{\boldsymbol{\Lambda}}_{nn}=\sigma\big(\boldsymbol{\theta}^{\top}\boldsymbol{x}_{n}\big)\big(1-\sigma(\boldsymbol{\theta}^{\top}\boldsymbol{x}_{n}\big)\big)$. We use Euclidean MAP, and here the negative Hessian coincides with the Fisher metric (Girolami and Calderhead,, 2011) (so both choices for $\operatorname{\boldsymbol{\Sigma}}$ are the same).
@@ 5.4 Neural network regression
Finally, we ran an experiment similar to the one by Bergamin et al., (2023), with data from Snelson and Ghahramani, (2005). The task is a 11D regression problem using an NN of size 11-1010-11 with $\tanh$ activation. The methods are implemented using Daxberger et al., (2021), with Fisher precision (the default option for full LA), $\hat{\operatorname{\boldsymbol{\theta}}}$ found by standard MAP training, and post hoc optimized prior precision and noise std. MSE and NLL are computed using $500$ samples.
@@
Table 3: NN regression results as $\text{mean},\text{std}]$. Bold font indicates the best method. $T$ indicates the average number of function evaluations for one sample while time indicates the average time for one sample. For all evaluation metrics smaller is better.
@@ 8.1 Theorem 1
The negative Hessian is therefore constant. For the second case, where the prior is uniform, the negative Hessian is 0 everywhere, thus is also constant.
@@ 8.3 Theorem 3
[…]
@@ 9 USING HAUSDORFF MAP AND FISHER PRECISION
[…]
@@ 11 BIAS OF RLA-B
Consider a positive $x$. It is clear that $\frac{1}{\sqrt{1+Dx^{2}}}$ decreases as $D$ increases. For smaller $D$, the expected velocity as calculated above may increase as $D$ increases, since $\operatorname{\mathbb{E}}\left[v_{0}\right]$ itself may increase. However, as $D$ approaches infinity, $\operatorname{\mathbb{E}}\left[\sqrt{D}v_{0}\right]$ approaches $\sqrt{D-1}$, $\operatorname{\mathbb{E}}\left[v_{0}\right]$ approaches 11 and $\operatorname{\mathbb{E}}\left[v\right]$ approaches 0.
@@ 14.2 Neural Networks (NN)
where $\boldsymbol{J}_{n}$ == $\nabla(\boldsymbol{\operatorname{\boldsymbol{\phi}}}_{n})_{1}\cdots\nabla(\boldsymbol{\operatorname{\boldsymbol{\phi}}}_{n})_{P}]^{\top}$ denotes the $P\times D$ Jacobian matrix of the NN at the $n^{th}$ input.
@@ 15.3 Finding the MAP estimate
For experiments other than neural network experiments, unless there is an analytical solution of the MAP estimate, we find the MAP estimate by collecting $20$ runs of BFGS optimization with random initialization using the implementation from Virtanen et al., (2020) which yield negative log-posterior values other than plus infinity or NaN, with maximum number of iterations $1e6$. We choose the run that yields the smallest negative log-posterior with numpy.argmin.
For neural network experiments, we find the MAP estimate using lr=1e2lr=1e-2, weight_decay=1e5weight\_decay=1e-5 for $20000$ epochs with the Adam (Kingma and Ba,, 2015) optimizer. The prior precision and noise std are optimized post hoc following standard procedure from Daxberger et al., (2021). During the optimization process, we fix the random seeds, in order to have consistent prior precision and noise std for NUTS.
@@ 15.5 ODE integration
For the exponential maps, other than neural network experiments, ODE integrations are carried out using Diffrax (Kidger,, 2021). Inspired by the default settings of the function solve_ivp in SciPy (Virtanen et al.,, 2020), we use Dormand-Prince’s $5(4)$ method (Dormand and Prince,, 1980). Diffrax reports the number of steps num_steps used by the integrator. In order to make the results approximately comparable to the number of function evaluations nfev reported by SciPy, we multiply num_steps by $6$, the number of function evaluations per step of Dormand-Prince’s $5(4)$ method. We use adaptive step size with initial step size chosen by the algorithm, with rtol=1e3rtol=1e-3 and atol=1e6atol=1e-6, matching the default option of SciPy. We set the maximum number of steps allowed to be $4096$ (approximately equivalent to $24576$ function evaluations), following the default configuration of Diffrax. For neural network experiments, following Bergamin et al., (2023), we use SciPy’s solve_ivp and use the default hyper parameters.
@@ 15.6 Calculating the Wasserstein distance
where $\operatorname{\boldsymbol{\gamma}}$ is the transportation plan, $\operatorname{\boldsymbol{M}}$ is the distance matrix induced by the Euclidean metric for the two sets of samples, and $\operatorname{\boldsymbol{a}}$ and $\operatorname{\boldsymbol{b}}$ are uniform numbers summing up to $1$ with the lengths the same as the numbers of respective samples. The maximum number of iterations is set to be $1e10$.
@@ 15.10 Neural networks
In Section 5.4, we performed experiments on NN regression tasks. Similar to Bergamin et al., (2023), we use the dataset from Snelson and Ghahramani, (2005), which has $200$ data points with labels, with the $x$ coordinates distributed between 0 and $6$.
@@ 16.5.2 Standardized data
In the neural network regression experiment from the main paper, we directly run the neural network on the original dataset from Snelson and Ghahramani, (2005). We later observed that using standardized data may lead to reduced running times for all methods while reducing possible numerical issues, and we report the results in Table 5 and Figure 14. For finding the MAP estimate, we set weight_decay=1e4weight\_decay=1e-4, with the other settings the same as for non-standardized data.
@@
Table 5: NN regression results with standardized data as $\text{mean},\text{std}]$. $T$ indicates the average number of function evaluations for one sample while time indicates the average time for one sample. Bold font indicates the best method. For all evaluation metrics smaller is better.